Calculate the pH of a buffer solution containing phenol and sodium phenoxide in a 1 to 3 ratio.
Answer #1

pKa of phenol=10

Here acid is phenol and conjugate base is sodium phenoxide.

Phenol concentration=[Acid]=C6H5OH=x M

Sodium phenoxide concentration =[Conjugate base]=C6H5ONa]=3x M

pH of a buffer is given by

Henderson - Hasselbalch equation of a buffer :

pH = pKa + log [conjugate base] / [acid]

pH=10 +log[3x]/[x]=10+log(3)=10+0.47=10.47

Thank you.